OpenStudy (hope210):

Help with algebra 1.

OpenStudy (helpppp):

Question?

OpenStudy (hope210):

evaluate: \[\sqrt[3]{343} +\frac{ 3 }{ 4 } \sqrt[3]{-8}\]

OpenStudy (helpppp):

@Rosie62122020

OpenStudy (hope210):

My options are \[-5\frac{ 1 }{ 2 }\] \[5\frac{ 1 }{ 2 }\] \[8\frac{ 1 }{ 2 }\]

OpenStudy (hope210):

@zepdrix

OpenStudy (helpppp):

use this. http://mathway.com